# Subgraph-Oriented Testing for Deep Learning Library

### v1

@@ Abstract
Deep Learning (DL) libraries, such as PyTorch, are widely used for building and deploying DL models on various hardware platforms. Meanwhile, they are found to contain bugs that lead to incorrect calculation results and cause issues like non-convergence training and inaccurate prediction of DL models. Thus, many efforts have been made to test DL libraries and reveal bugs. However, existing DL library testing methods manifest limitations: model-level testing methods cause complexity in fault localization. Meanwhile, API-level testing methods often generate invalid inputs or primarily focus on extreme inputs that lead to crash failures; they also ignore testing realistic API interactions. These limitations may lead to missing detection of bugs, even in the frequently used APIs. To address these limitations, we propose SORT (Subgraph-Oriented Realistic Testing) to differential test DL libraries on different hardware platforms. SORT takes popular API interaction patterns, represented as frequent subgraphs of model computation graphs, as test subjects. In this way, it introduces realistic API interaction sequences while maintaining efficiency in locating faulty APIs for observed errors. Besides, SORT prepares test inputs by referring to extensive features of runtime inputs for each API in executing real-life benchmark data. The generated inputs are expected to better simulate such valid real inputs and reveal bugs more likely to happen in real-life usage. Evaluation on 728 frequent subgraphs of 49 popular PyTorch models demonstrates that SORT achieves a 100% valid input generation rate, detects more precision bugs than existing methods, and reveals interaction-related bugs missed by single-API testing. 18 precision bugs in PyTorch are identified.
@@ I Introduction
To avoid the shortcomings of the bulky model-level testing, many API-level DL library testing methods have emerged recently [14, 16, 17, 18, 19, 20]. Such methods test APIs in a DL library individually. Specifically, they generate test inputs for one single API and check the correctness of corresponding outputs. Such testing methods address the limitations of model-level methods and become a popular paradigm. However, they also manifest three significant problems.
@@
On 728 subgraphs extracted from 49 popular DL models implemented in PyTorch, we did differential testing across CPU and GPU with SORT to evaluate its effectiveness. Results show that SORT can effectively reveal precision errors with valid inputs. Specifically, SORT shows a 100% valid rate of generated inputs, outperforming all baselines. Besides, SORT detects more precision errors than baselines. It is also found that many reported bugs initially appear as small precision losses, which cannot be detected without considering API interactions. In addition, we conduct a survey to collect developers’ thoughts about DL library testing. The feedback confirms the helpfulness of taking frequent graphs as testing subjects in introducing meaningful API interactions while maintaining fault localization efficiency, as well as the significance of the revealed precision bugs signified by API interactions.
The primary contributions of this paper can be summarized as follows:
  • We propose to test DL libraries at a new granularity, i.e., taking frequent subgraphs of model computation graphs as test subjects. Subgraph-level testing introduces more meaningful API interactions to expose the impactful output differences more likely to happen in the real-life usage of DL libraries while maintaining efficiency in locating the root cause of the observed differences.
@@
  • We revealed 18 precision bugs of PyTorch.
II Preliminary
@@ II-A1 Lack of Valid Inputs
However, existing methods report a low valid rate of generated inputs. For example, DocTer [16] and ACETest [21] report valid rates of 12.38% and 43.66%. DocTer’s low valid rate can be attributed to incomplete documentation [16] and ACETest’s relatively low ratio of valid inputs is due to limited information extracted from the code [21]. We also reproduced the other popular existing testing methods that do not report the ratio of valid inputs, i.e., FreeFuzz [14], TitanFuzz [18], and DeepREL [17], and found that their average ratio of valid inputs is 16.01%, 41.54%, and 81.34%, respectively. Overall, the ratio of valid inputs generated by current methods is unsatisfactory, mainly due to incomplete or unreliable information to guide generation. The detailed reasons will be discussed in Section V-A. The unsatisfactory ratio of valid inputs may result in inadequate testing of APIs’ calculation function [21].
II-A2 Lacking Detection of Precision Errors
Even if existing methods can generate a few test inputs that pass APIs’ validity check, most such inputs are corner cases that lead to crash failures, e.g., tensors with extreme values or unusual types. These crash failures tend to happen less frequently in real-life model deploying scenarios as suggested by developers. As shown in Fig. 1, the PyTorch developers considered many crash issues reported by TitanFuzz (highlighted in a red box) as “not real user issues” [^1] https://github.com/pytorch/pytorch/issues/88148. Meanwhile, other types of errors need to be exposed by non-corner cases. Precision errors are typical of such errors. Precision errors will lead to problematic calculation results with precision loss compared with the correct result. Such errors are considered a central concern in high-performance computing since significant precision errors will lead to inaccurate calculation results [22] and damage supercomputers and handheld electronics systems [27].
However, current methods often overlook generating non-corner cases and seldom identify precision errors. For example, FreeFuzz, DeepREL, DocTer, TitanFuzz, and ACETest have identified a total of 143 issues of PyTorch; but only four of them are precision-related: FreeFuzz found one, TitanFuzz reported three, and DeepREL, DocTer, and ACETest reported no precision issues. Although Predoo and NablaFuzz aim to identify precision errors, they only apply to few specific APIs. Predoo only applies to seven APIs with known susceptibility to precision issues and NablaFuzz only focuses on automatic differentiation components for gradient calculation.
@@ II-A2 Lacking Detection of Precision Errors
II-A3 Lack of Realistic Interactions
@@
However, traditional API-level testing methods test each API independently and cannot detect such errors via API interactions. Recently, DeepREL [17] and TitanFuzz [18] artificially pair APIs to make up API interactions, but errors found by these artificial interactions are usually considered as not occurring in real model-executing scenarios. For instance, developers regarded issue #87960 in Fig. 1, which is caused by an API interaction made up by TitanFuzz, as “not a real user issue”. These suggest the need to consider API interactions in testing, particularly the realistic interactions such that the errors more likely to affect reliability in real model-deploying scenarios can be better exposed.
@@
output = f.layer_norm(p0, p1, p2, p3, p4)
@@
Figure 2: Precision Differences of Single API Testing
@@
# Precision Differences: 0.17188
Figure 3: Precision Differences of Real Interaction Testing
II-B A New Granularity for DL Library Testing
To address the limitations discussed in Section II-A, we introduce a new granularity of DL library testing subjects: frequent subgraphs. Specifically, a deep learning model can be represented as a computational graph, where each node represents an API used in calculation and each edge represents the data flow between APIs. Frequent subgraphs refer to the subgraphs frequently appearing in the computation graphs of a series of DL models. For example, Figs. 4 and  5 show part of the definition code of ResNet and DenseNet models, respectively. Accordingly, we can obtain their computation graphs as shown in Fig. 6. Then, we can identify frequent subgraphs with a frequency of two from the two models as highlighted with the frame, where the output of conv2d is input to batch_norm, and the output from batch_norm is subsequently input to relu. Such frequent subgraphs represent popular real interaction patterns of DL library APIs.
@@ II-B A New Granularity for DL Library Testing
Figure 4: Part of the Definition Code of the ResNet Model
@@
Figure 5: Part of the Definition Code of the DenseNet Model
Figure 6: ResNet and DenseNet Model Graph
Testing DL libraries at the granularity of frequent subgraphs manifests benefits. Specifically, different from the individual APIs in traditional API-level testing, frequent subgraphs introduce API interactions to magnify the small differences likely significantly impacting the calculation result. Moreover, the frequent subgraphs identified from real DL models introduce meaningful API interactions rather than artificially made-up combinations whose significance is uncertain. As a reminder, using whole DL models to test DL libraries also introduces realistic API interactions, but DL models often include many APIs and are thereby hard to localize the problematic API [14]. Meanwhile, frequent subgraphs are smaller than the whole DL model, while they include common API interactions in popular DL models, which allows prioritizing the validation of these widely-used interactions. According to the data we collected, a DL model has an average of 339 APIs, with a minimum of 29 APIs and a maximum of 1143 APIs; meanwhile, subgraphs have only six APIs on average, with a minimum of two APIs and a maximum of seven APIs. Consequently, testing frequent subgraphs maintains a better efficiency of bug localization.
@@ III Methodology
In this work, we propose a DL library testing method named SORT. SORT tackles the limitations of existing methods by differential testing frequent subgraphs’ behaviors with test inputs generated based on real-life inputs’ features as elaborated in Section II-B. As shown in Fig. 7, SORT includes four modules. The first module extracts the frequent subgraphs from open-source DL models as the test subjects; the second module collects runtime input data features of open-source DL models; the third module generates test inputs for frequent subgraphs by referring to the runtime input data; the last module performs differential testing over frequent subgraphs with the generated inputs across hardware platforms to expose problematic different behaviors, e.g., precision errors. We elaborate on these modules in the following sections.
III-A Frequent Subgraphs Extraction
@@
Once the complete computation graph for each popular model is built, we apply a frequent subgraph extraction algorithm, gSpan [29], to identify the frequent subgraphs. gSpan iterates through a set of graphs and conducts a depth-first search to discover and expand subgraphs that meet a minimum frequency threshold, which represents the minimum number of times a subgraph must appear across the model graphs to be considered “frequent”. For example, if we use a frequency threshold of 2, then we can identify a frequent subgraph made up of APIs conv2d, batch_norm, and relu based on the ResNet and DeseNet models as shown in Fig. 6.
III-B Runtime Input Data Features Collection
In this section, we introduce how we record the features of runtime data of DL library APIs. Such information is later used to guide the generation of test inputs for frequent subgraphs, which will be introduced in Section III-C.
To collect the features of API runtime data, we instrument the source code of each DL library API and run real DL models with real datasets. Specifically, we hook each API call to record input information during each execution. For the tensor parameters, we record information about the type, shape, and value range. For the non-tensor parameters, the values are recorded. When an API is invoked once in a model with a real input for the model, we will collect a piece of input information. For example, Fig. 8 shows a piece of information we collect for API conv2d. It suggests that for the parameter input of API conv2d, we can generate a tensor with a type of float32 and a shape of [16, 3, 224, 224] and values ranging from -102.9072 to 152.3829 as a reasonable test input; for the parameter stride, we can generate an integer test input of 1; etc. In this way, we can collect multiple pieces of input information for each API after running several models invoking the API with diverse model inputs. The collected information guides the later generation of test inputs for the corresponding API.
@@ III-B Runtime Input Data Features Collection
Figure 8: A Piece of Input Information for conv2d
@@ III-C Test Input Generation
With the collected information, we generate test inputs for each frequent subgraph. To systematically generate the input values for parameters, we categorize parameters into three types: parameters of entry API, dependent parameters of non-entry API, and non-dependent parameters of non-entry API. In the following, we explain these parameter types and our method to generate values for them.
@@
Figure 9: A Piece of Input Information for batch_norm
@@ III-D Test Case Execution and Result Verification
The last step of SORT is to perform differential testing over each frequent graph using the generated test inputs. We follow existing works [23, 14, 18] to run the test cases on two different hardware platforms, e.g. CPU and GPU, and compare their outputs. If the difference between the outputs obtained on different platforms is within an acceptable tolerance threshold, we consider there is no error; otherwise, potential bugs are reported. This criterion is originally suggested by the PyTorch community [^3] https://discuss.pytorch.org/t/precision-difference-between-gpu-and-cpu/26969,https://discuss.pytorch.org/t/computation-precision-differs-a-lot-between-gpu-mode-and-cpu-mode/23112, where developers introduce the precision issues due to the inherent differences between CPU and GPU and highlight the importance of exposing them.
@@ IV-A Research Questions
RQ3: Can interactions contribute to detecting more bugs? In this RQ, we compare the effect of testing with and without considering API interaction to understand if the API interaction patterns reflected by frequent subgraphs contribute to the bug-detection ability of SORT.
@@ IV-D Metrics
Categories of Bug Types. We categorized the revealed bugs based on the symptoms [30, 31, 32, 9]. Specifically, we consider four categories: crash bugs, NaN bugs, precision bugs, and performance bugs. Crash bugs refer to instances where the program exits unexpectedly in execution with an error. NaN bugs and precision bugs occur when the API returns a result of NaN or other value that is inconsistent with the expected one, respectively. Performance degradation bugs occur when the program exhibits poor performance or overuses computational resources, such as taking an excessive amount of time to produce results [31].
@@ IV-E Implementation
Subgraph Extraction. We collected models implemented in PyTorch and can be successfully run on both CPU and GPU from OpenMMLab model zoo, which results in 49 DL models. According to our preliminary experiments, when we set the threshold to be 5, we can collect the top 20% frequently appearing subgraphs (API interaction patterns). In this work, we use these frequent subgraphs (728 in total) as our test subjects. Considering the popularity of these API interactions, they are expected to have been extensively tested. We try to reveal the bugs in these APIs missed by the existing methods. There are 15 APIs in these frequent subgraphs: flatten, __mul__, div, softmax, adaptive_avg_pool2d, matmul, max_pool2d, batch_norm, dropout, relu, conv2d, gelu, linear, layer_norm, and __add__. Our evaluation will focus on testing these essential APIs in popular API usage patterns. As a reminder, our approach is not restricted to these 15 APIs. This setup is made to evaluate if our method can help to reveal the errors that reside in very popular APIs yet are missed by existing methods. By using a lower frequency threshold, SORT can test a broader range of APIs.
Test Execution. In our experiments, we run the extracted frequent subgraphs on both CPU and GPU and required the outputs of these APIs in subgraphs to adhere that any element of the output tensor produced by these executions should satisfy the following condition:
@@
By setting this criterion following [14, 33], we expect to identify potential precision bugs.

### v2

# Subgraph-Oriented Testing for Deep Learning Libraries

@@ Abstract
Deep Learning (DL) libraries, such as PyTorch, are widely used for building and deploying DL models on various hardware platforms. Meanwhile, they are found to contain bugs that lead to incorrect calculation results and cause issues like non-convergence training and inaccurate prediction of DL models. Thus, many efforts have been made to test DL libraries and reveal bugs. However, existing DL library testing methods manifest limitations: model-level testing methods cause complexity in fault localization. Meanwhile, API-level testing methods often generate invalid inputs or primarily focus on extreme inputs that lead to crash failures; they also ignore testing realistic API interactions. These limitations may lead to missing detection of bugs, even in the frequently used APIs. To address these limitations, we propose SORT (Subgraph-Oriented Realistic Testing) to differential test DL libraries on different hardware platforms. SORT takes popular API interaction patterns, represented as frequent subgraphs of model computation graphs, as test subjects. In this way, it introduces realistic API interaction sequences while maintaining efficiency in locating faulty APIs for observed errors. Besides, SORT prepares test inputs by referring to extensive features of runtime inputs for each API in executing real-life benchmark data. The generated inputs are expected to better simulate such valid real inputs and reveal bugs that are more likely to happen in real-life usage. Evaluation on 728 frequent subgraphs of 49 popular PyTorch models demonstrates that SORT achieves a 100% valid input generation rate, detects more precision bugs than existing methods, and reveals interaction-related bugs missed by single-API testing. 18 precision bugs in PyTorch are identified and reported to PyTorch developers.
@@ I Introduction
To avoid the shortcomings of the bulky model-level testing, many API-level DL library testing methods have emerged recently [14, 16, 17, 18, 19, 20, 21]. Such methods test APIs in a DL library individually. Specifically, they generate test inputs for one single API and check the correctness of corresponding outputs. Such testing methods address the limitations of model-level methods and have become a popular paradigm. However, they also manifest three significant problems.
@@
On 728 subgraphs extracted from 49 popular DL models implemented in PyTorch, we performed differential testing across CPU and GPU using SORT to evaluate its effectiveness. Results show that SORT can effectively reveal precision errors with valid inputs. It shows a 100% validity rate of generated inputs, outperforming all baselines. Furthermore, SORT detects more precision errors than baselines. It is also found that many reported bugs initially appear as small precision losses, which cannot be detected without considering API interactions. In addition, we conduct a survey to collect developers’ thoughts about DL library testing. The feedback confirms the helpfulness of taking frequent graphs as testing subjects in introducing meaningful API interactions while maintaining fault localization efficiency, as well as the significance of the revealed precision bugs signified by API interactions.
The contributions of this paper are summarized as follows:
  • We propose to test DL libraries with a new granularity of test subjects, i.e., frequent subgraphs of model computation graphs. Testing with frequent subgraphs introduces more meaningful API interactions to expose the impactful output differences that are more likely to happen in the real-life usage of DL libraries while maintaining efficiency in locating the faulty APIs causing the observed differences.
@@
  • We identified 18 precision bugs of PyTorch.
We release the artifact of this paper including the replication package and additional materials online at [27].
II Motivation and Preliminary Study
@@ II-A1 Lack of Valid Inputs
However, existing methods report a low valid rate of generated inputs. For example, DocTer [16] and ACETest [19] report valid rates of 12.38% and 43.66%, respectively. DocTer’s low valid rate can be attributed to incomplete documentation [16] and ACETest’s relatively low ratio of valid inputs is due to limited information extracted from the code [19]. We also reproduced the other popular existing testing methods that do not report the ratio of valid inputs, i.e., FreeFuzz [14], TitanFuzz [18], and DeepREL [17], and found that their average ratio of valid inputs is 16.01%, 41.54%, and 81.34%, respectively. Overall, the ratio of valid inputs generated by current methods is unsatisfactory, mainly due to incomplete or unreliable information to guide generation. The detailed reasons will be discussed in Section V-A. The unsatisfactory ratio of valid inputs may result in inadequate testing of APIs’ calculation function [19].
II-A2 Missing Reporting Precision Errors
Even if existing methods can generate a few test inputs that pass APIs’ validity check, most such inputs are corner cases that lead to crash failures, e.g., tensors with extreme values or unusual types. These crash failures tend to happen less frequently in real-life model deploying scenarios as suggested by developers. As shown in Fig. 1, the PyTorch developers considered many crash issues reported by TitanFuzz (highlighted in a red box in the figure) as “not real user issues” [^1] https://github.com/pytorch/pytorch/issues/88148. Meanwhile, other types of errors need to be exposed by non-corner cases. Precision errors are typical of such errors, which manifest as problem-inducing value differences (i.e., precision losses) between the correct calculation result and the actual results returned on certain computation hardware. They arise from the differences of unique hardware in handling limitations of finite precision and rounding errors inherent in floating-point representations [20, 28]. Such errors are considered a central concern in high-performance computing since significant precision errors will lead to inaccurate calculation results [22] and damage supercomputers and handheld electronics systems [28].
However, current methods often overlook generating non-corner cases and seldom identify precision errors. For example, FreeFuzz, DeepREL, DocTer, TitanFuzz, and ACETest have identified a total of 143 issues of PyTorch; but only four of them are precision-related: FreeFuzz found one, TitanFuzz reported three, and DeepREL, DocTer, and ACETest reported no precision issues. Although Predoo and NablaFuzz aim to identify precision errors, they only apply to a few specific APIs. Predoo only applies to seven APIs with known susceptibility to precision issues and NablaFuzz only focuses on automatic differentiation components for gradient calculation.
@@ II-A2 Missing Reporting Precision Errors
II-A3 Lack of Realistic API Interactions
@@
However, traditional API-level testing methods test each API independently and cannot detect such errors via API interactions. Recently, TitanFuzz [18] artificially pairs APIs to make up API interactions, but errors found by these artificial interactions are usually considered as not occurring in real model-executing scenarios. For instance, developers regarded issue #87960 in Fig. 1, which is caused by an API interaction made up by TitanFuzz, as “not a real user issue”. These suggest the need to consider API interactions in testing, particularly the realistic interactions such that the errors more likely to affect reliability in real model-deploying scenarios can be better exposed.
@@
output_ln = f.layer_norm(p0, p1, p2, p3, p4)
@@
Figure 2: Example Precision Differences in Testing Single API
@@
# Accumulated Precision Differences: 0.17188
Figure 3: Example Precision Differences in Testing with API Interactions
II-B Frequent Subgraph-Level DL Library Testing
To address the limitations discussed in Section II-A, we introduce a new granularity of DL library testing subjects: frequent subgraphs. Specifically, a deep learning model can be represented as a computational graph, where each node represents an API used in calculation and each edge represents the data flow between APIs. Frequent subgraphs refer to the subgraphs frequently appearing in the computation graphs of a series of DL models. For example, Figs. 4 and  5 show part of the definition code of ResNet [30] and DenseNet [31] models, respectively. Accordingly, we can obtain their computation graphs as shown in Fig. 6. Then, we can identify frequent subgraphs with a frequency of two from the two models as highlighted with the frame, where the output of conv2d is input to batch_norm, and the output from batch_norm is subsequently input to relu. Such frequent subgraphs represent popular real interaction patterns of DL library APIs.
@@ II-B Frequent Subgraph-Level DL Library Testing
Figure 4: Part of the Code Defining the ResNet Model
@@
Figure 5: Part of the Code Defining the DenseNet Model
Figure 6: Computation Graphs of ResNet and DenseNet Models
Testing DL libraries at the frequent subgraphs level manifests benefits. Specifically, different from the individual APIs in traditional API-level testing, frequent subgraphs introduce API interactions to magnify the small differences likely significantly impacting the calculation result. Moreover, the frequent subgraphs identified from real DL models introduce meaningful API interactions rather than artificially made-up combinations whose significance is uncertain. As a reminder, using whole DL models to test DL libraries also introduces realistic API interactions, but DL models often include many APIs and are thereby hard to localize the problematic API [14]. Meanwhile, frequent subgraphs are smaller than the whole DL model, while they include common API interactions in popular DL models, which allows prioritizing the validation of these widely-used interactions. According to the data we collected, a DL model has an average of 339 APIs, with a minimum of 29 APIs and a maximum of 1,143 APIs; meanwhile, subgraphs have only six APIs on average, with a minimum of two APIs and a maximum of seven APIs. Consequently, testing frequent subgraphs maintains a better efficiency of bug localization.
@@ III Methodology
In this work, we propose a DL library testing method named SORT. SORT tackles the limitations of existing methods by differential testing frequent subgraphs’ behaviors with test inputs generated based on real-life inputs’ features as elaborated in Section II-B. As shown in Fig. 7, SORT includes four modules. The first module extracts the frequent subgraphs from open-source DL models as the test subjects; the second module collects runtime input data features of open-source DL models; the third module generates test inputs for frequent subgraphs by referring to the runtime input data; the last module performs differential testing over frequent subgraphs with the generated inputs across hardware platforms to expose problematic different behaviors, e.g., precision errors. We elaborate on these modules in the following subsections.
III-A Frequent Subgraph Extraction
@@
Once the complete computation graph for each collected model is built, we apply a frequent subgraph extraction algorithm, gSpan [32], to identify the frequent subgraphs. gSpan is one of the most popular solutions for mining frequent subgraphs [33]. It iterates through a set of graphs and conducts a depth-first search to discover and expand subgraphs that meet a minimum frequency threshold, representing the minimum times a subgraph must appear across the model graphs to be considered “frequent”. For example, if we use a frequency threshold of 2, then we can identify a frequent subgraph made up of APIs conv2d, batch_norm, and relu based on the ResNet and DeseNet models as shown in Fig. 6.
III-B Runtime Input Feature Information Collection
In this subsection, we introduce how we record the features of runtime data of DL library APIs. Such information is later used to guide the generation of test inputs for frequent subgraphs, which will be introduced in Section III-C.
To collect the features of API runtime data, we instrument the source code of each DL library API and run real DL models with real datasets. Specifically, we hook each API call to record input feature information during each execution. For the tensor parameters, we record feature information about the type, shape, and value range. For the non-tensor parameters, the values are recorded. When an API is invoked once in a model with a real input for the model, we will collect a piece of input feature information. For example, Fig. 8 shows a piece of input feature information we collect for API conv2d. It suggests that for the parameter input of API conv2d, we can generate a tensor with a type of float32 and a shape of [16, 3, 224, 224] and values ranging from -102.9072 to 152.3829 as a reasonable test input; for the parameter stride, we can generate an integer test input of 1. In this way, we can collect multiple pieces of input feature information for each API after running several models invoking the API with diverse model inputs. The collected feature information guides the later generation of test inputs for the corresponding API.
@@ III-B Runtime Input Feature Information Collection
Figure 8: Recorded Input Feature Information of conv2d
@@ III-C Test Input Generation
With the collected feature information, we generate test inputs for each frequent subgraph. To systematically generate the input values for parameters, we categorize parameters into three types: parameters of entry API, dependent parameters of non-entry API, and non-dependent parameters of non-entry API. In the following, we explain these parameter types and our method to generate values for them.
@@
Figure 9: Recorded Input Feature Information of batch_norm
@@ III-D Test Case Execution and Result Verification
The last step of SORT is to perform differential testing over each frequent graph using the generated test inputs. We follow existing works [20, 14, 18] to run the test cases on two different hardware platforms, e.g. CPU and GPU, and compare their outputs. If the difference between the outputs obtained on different platforms is within an acceptable tolerance threshold, we consider there is no error; otherwise, potential bugs are reported. This criterion is originally suggested by the PyTorch community [^3] https://discuss.pytorch.org/t/precision-difference-between-gpu-and-cpu/26969,https://discuss.pytorch.org/t/computation-precision-differs-a-lot-between-gpu-mode-and-cpu-mode/23112, where users introduce the precision issues due to the inherent differences between CPU and GPU and developers highlight the importance of exposing them.
@@ IV-A Research Questions
RQ3: Can API interactions contribute to detecting more bugs? In this RQ, we compare the effect of testing with and without considering API interaction to understand if the API interaction patterns reflected by frequent subgraphs contribute to the bug-detection ability of SORT.
@@ IV-D Metrics
Categories of Bug Types. We categorized the revealed bugs based on the symptoms [34, 35, 36, 9]. Specifically, we consider four categories: crash bugs, NaN bugs, precision bugs, and performance degradation bugs. Crash bugs refer to instances where the program exits unexpectedly in execution with an error. NaN bugs and precision bugs occur when the API returns a result of NaN or other value that is inconsistent with the expected one, respectively. Performance degradation bugs occur when the program exhibits poor performance or overuses computational resources, such as taking an excessive amount of time to produce results [35].
@@ IV-E Implementation
Extraction of Frequent Subgraphs. We collected models implemented in PyTorch and can be successfully run on both CPU and GPU from OpenMMLab model zoo, which results in 49 DL models. According to our preliminary experiments, when we set the threshold to be 5, we can collect the top 20% frequently appearing subgraphs (API interaction patterns). In this work, we use these frequent subgraphs (728 in total) as our test subjects. Considering the popularity of these API interactions, they are expected to have been extensively tested. We try to reveal the bugs in these APIs missed by the existing methods. There are 15 APIs in these frequent subgraphs: flatten, __mul__, div, softmax, adaptive_avg_pool2d, matmul, max_pool2d, batch_norm, dropout, relu, conv2d, gelu, linear, layer_norm, and __add__. Our evaluation will focus on testing these essential APIs in popular API usage patterns. As a reminder, our approach is not restricted to these 15 APIs. This setup evaluates whether our method can help reveal the errors that reside in very popular APIs yet are missed by existing methods. By using a lower frequency threshold, SORT can test a broader range of APIs.
Collection of Runtime Input Features. To collect runtime inputs for each API, we executed the 49 popular DL models from OpenMMLab using the datasets recommended by OpenMMLab for each model. For example, we ran the Conformer [37] model on the ImageNet [38] dataset and the ResNet [30] model on the CIFAR [39] dataset to collect runtime input features for the APIs invoked when running the two models.
Test Execution. In our experiments, we ran the extracted frequent subgraphs on both CPU and GPU and required the outputs of these APIs in subgraphs to adhere that any element of the output tensor produced by these executions should satisfy the following condition:
@@
where a and b refer to the float-point values at each index in the output tensors in the CPU and GPU executions, respectively. By setting this criterion following [14, 40], we expect to identify potential precision bugs.